OpenStudy (anonymous):

When a driver presses the brake pedal, his car stops with an acceleration of -0.5 m/s2. How far will the car travel while coming to a complete stop if its initial speed was 8 m/s? A. 16 m B. 32 m C. 64 m D. 8 m

OpenStudy (cybershadow):

see we have been given initial speed = u = 8 m/s and acceleration a = -0.5 m/s^2 and we need to find how much distance will the car travel before comin to rest i.e. final speed v =0 so we know that v^2-u^2 = 2as we need to find s 0^2 - 8^2 = 2 * -0.5 * S solvint it we get S = -64/-1 S = 64 m so Option C

OpenStudy (anonymous):

Alternatively, find how long the cars slows down, t = v/a = 8/.5 = 16 sec. Next use s = v t - (1/2) a t^2 = 8*16 - (0.25) (16)^2 = 64, agreeing, as they should.